OpenStudy (anonymous):

2/3a x 2/3a^2 please simplify i got 4/3a^2 i just want someoneto check it.

OpenStudy (anonymous):

seems right to me.

OpenStudy (jdoe0001):

ahemm, \(3a = 3\times a^1 \implies 3a^1 \times 3a^2 =\ ?\)

OpenStudy (anonymous):

idk what u mean?

OpenStudy (jdoe0001):

is this what you have \( \cfrac{2}{3a}\times \cfrac{2}{3a^2} \ ? \)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so, what would you get for \(\large 3a = 3\times a^1 \implies 3a^1 \times 3a^2 =\ ? \)

OpenStudy (anonymous):

3a^2

OpenStudy (jdoe0001):

if you checked the picture above, for the exponential properties :|

OpenStudy (anonymous):

is my answer right or not?

OpenStudy (jdoe0001):

well, \(2\times2=4\), so, the numerator is ok

OpenStudy (anonymous):

so ur saying my denom is wrong....

OpenStudy (jdoe0001):

now, if \( 3a^1 \times 3a^2 = 3a^2\) that means that \(\boxed{3a^1 \times 3a^2} = \boxed{3a^2}\) are they if using the exponent properties?

OpenStudy (anonymous):

no?

OpenStudy (jdoe0001):

dunno, look at the picture, the 1st of the properties, also you're skipping on the coefficient

OpenStudy (anonymous):

so its 3a^1 x 3a^@ is 3a^3?

OpenStudy (anonymous):

3a^2

OpenStudy (jdoe0001):

\(a^3\) is ok, you're still skipping on the coefficients

OpenStudy (anonymous):

what are they then?

OpenStudy (anonymous):

do u want me to multiply the 3?

OpenStudy (anonymous):

like with the 3a^1 and 3a^2?

OpenStudy (jdoe0001):

well, yes :|

OpenStudy (anonymous):

please explain what im doing wrong here jdoe?

OpenStudy (jdoe0001):

$$ 3a^1 \times 3a^2 = (\color{blue}{3\times3})(a^{\color{blue}{1+2}}) $$